OpenStudy (anonymous):

What is the common ratio of the following sequence? -5, -20, -80, ...

OpenStudy (anonymous):

Sn=4*Sn-1, where S1=-5

OpenStudy (anonymous):

what is -20/-5 and likewise -80/-20?? There's your answer :)

OpenStudy (anonymous):

options are 4 , -4 and 1/4

OpenStudy (anonymous):

so what is -20/-5?? use a calculator if you must. But that is the answer.

OpenStudy (anonymous):

-20 ÷ -5 -80 ÷ -20

OpenStudy (anonymous):

Whats the comman ratio between -5, -20 ,-80

OpenStudy (anonymous):

I'm trying to get you to solve it...I'm just helping you -20 ÷ -5 = x -80 ÷ -20 = x x is going to be your answer

OpenStudy (anonymous):

you just apply each possible answer to the sequence and check the result as @rebeccaskell94 proposes...

OpenStudy (anonymous):

i found the answer its 4

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

thank